# The exact convex roof for GHZ-W mixtures for three qubits and beyond

### v1

@@ I Introduction
For the remaining points in the Bloch sphere there is only a single visible state from the zero polytope left which leads to $(1,1)$ decomposition being optimal there. They are depicted in Fig. 7 by green lines and red points.
@@
Conclusions – The mixture of GHZ and W state has been elaborated for all the Bloch sphere, obtaining a pattern on the surface that is characteristic for these two states in the mixture. The finding is based on the zero-state locking and that $(n_{z},n_{e})$ decompositions for $n_{e}>1$ are not optimal for the present case. We have checked this numerically but there is no general proof for that finding. For the zero-state locking it is essential to look at a certain roots of the tangle, which is the square root in the case of the threetangle. The Bloch sphere is divided into the zero polytope and four polytopes onto each faces of the zero polytope with corresponding four pure states. The faces of two neighboring such polytopes are strictly planar. There are pure states that belong to $(2,1)$ decompositions moving on three azimuthal grand circles and three circles which possess a small distance $0.0711148$ to the Bloch sphere center. Their normal vector is $(0.57589\cos\phi_{i},0,-0.81753\sin\phi_{i})$, with $\phi_{i}=i2\pi/3$. The remaining decompositions to fill the Bloch sphere are of the type $(1,1)$. This distribution of the optimal decomposition is natural and inter-connects the polytopes of various dimensions. For SL invariant tangles there is a transformation rule that applies in particular to the optimal decomposition statesViehmann et al. (2012). Applying these rules yields the solution for the convex-roof for these cases. For the general case the characteristic features of the convex roof will modify, but the structure will stay the same, as reported also in ref. Neveling and Osterloh (2024). It would be interesting to what extent the findings are universal. The paths of the pure entangled state of $(2,1)$ decompositions here moves on circles, but this does not need to be this way. Future investigation, clarifying this question, will be needed. The analysis is based on two essential pillars: the zero-state locking and that every optimal decomposition would be of the type $(n_{z},1)$. It would be crucial to either prove or disprove these conjectures in future work.
This presumably clarifies how optimal decompositions of rank two mixtures behave. It will be very intriguing to understand their behavior for higher ranks.
@@
Appendix A General properties of optimal decompositions for real states of rank two
@@
$\lambda=\frac{\mu_{1}}{\mu_{0}}=\frac{\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (6)
and for the mainly interesting case of $\phi_{1}=0$ and $p$ being the probability in the density matrix in consideration, we obtain as a solution to c)
@@
Figure 8: States within the Bloch sphere. For a given line $\overline{x_{1}x_{2}}$ all triangles $z_{0}x_{1}z_{1}$, $z_{0}x_{2}z_{2}$, and $z_{0}x_{2}^{\prime}z_{2}^{\prime}$, are similar. Ratios of their length are determined by the theorem of intersecting lines.
@@
A.1 $(1,1)$-decompositions
Here, we collect the formulae obtained for $(1,1)$-decompositions of a pure state at z-coordinate $12p_{1}-$ 1, where $p_{1}$ corresponds to its probability of the two states $|\psi_{i}\rangle$; $i\in\{N,S\}$. This corresponds to two pure states in the Fig. 8. We find
@@ A.1 $(1,1)$-decompositions
Next we want to examine whether for a given $(1,1)$ decomposition with $p_{1}$, $p_{2}$, and $p$ it is convenient to split the pure state into two, having a $(1,2)$ decomposition. It can be seen that the ratios $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ are equal with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}_{2}}^{2}}$ (see Fig. 8). For the new $p^{\prime}_{2}$ of this new state at phase $\phi_{2}$ the calculation gives
@@
A.2 $(2,2)$-decompositions
In the following we will assume that two complex conjugated pairs, corresponding to a superposition of two states at $p_{i}$ and angles ±$\pm\phi_{i}$, $i\in\{1,2\}$, such that both states come to lie inside the Bloch sphere projection. We find
@@ A.2 $(2,2)$-decompositions
Next, for the evaluation of $(2,2)$ from $(2,1)$ decompositions for given $p_{1}$, $\phi_{1}$, and $p$ the ratios of the following quantities are kept fixed (see Fig. 8): $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}}_{2}^{2}}$. These have to be substituted in $m_{2}(p_{1},\phi_{1};p^{\prime}_{2},\phi^{\prime}_{2})$ where $x^{\prime}_{2}=2\sqrt{p^{\prime}_{2}(1-p^{\prime}_{2})}\cos\phi^{\prime}_{2}$.

### v2

@@ I Introduction
For the remaining parts of the Bloch sphere there is only a single visible state from the zero polytope left and leads to $(1,1)$ decomposition being optimal. They are depicted in Fig. 1 by green lines and red points.
Conclusions – The mixture of GHZ and W state has been elaborated for all the Bloch sphere, obtaining a pattern on the surface that is characteristic for these two states in the mixture. We give a proof of the zero-state locking where the finding is based on. We also derive an inequality which decides whether $(n_{z},n_{e})$ decompositions for $n_{e}>1$ are optimal and find them not optimal for the present case. For the zero-state locking it is essential to look at a certain roots which render the tangle to scale linear with the density matrix. This is the square root in case of the threetangle.
@@
The Bloch sphere is divided into the zero polytope and four polytopes onto each faces of the zero polytope with corresponding four pure states. The faces of two neighboring such polytopes are strictly planar. There are pure states that belong to $(2,1)$ decompositions moving on three azimuthal grand circles and three circles which possess a small distance $0.0711148$ to the Bloch sphere center. Their normal vector is $(0.57589\cos\phi_{i},0,-0.81753\sin\phi_{i})$, with $\phi_{i}=i2\pi/3$. The remaining decompositions to fill the Bloch sphere are of the type $(1,1)$. This distribution of the optimal decomposition inter-connects the polytopes of various dimensions. For SL invariant tangles there is a transformation rule that applies in particular to the optimal decomposition statesViehmann et al. (2012). Applying these rules yields the solution for the convex-roof for the transformed density matrix. For the general case the characteristic features of the convex roof will modify, but the structure will stay the same, as reported also in ref. Neveling and Osterloh (2024). It would be interesting to what extent the findings are universal. The paths of the pure entangled state of $(2,1)$ decompositions here moves on circles, but this does not need to be general. It is intriguing to look for examples where the inequality points towards $(0,2)$ decompositions in parts of the Bloch sphere. Future investigations, clarifying these questions, will be needed. This work enters the way how optimal decompositions of rank two mixtures behave. It will be a challenge to understand their behavior for higher ranks. This can also become relevant for various quantum technologies. Future work should establish the specific role of genuine SL-invariant entanglement in this respect.
@@
Appendix A Detecting genuine three-partite entanglement via the threetangle
We will consider $\sqrt{|\tau_{3}|}$ as entanglement measure, where the threetangle $|\tau_{3}|$ has been defined asCoffman et al. (2000) (see also in Refs. Wong and Christensen (2001); Verstraete et al. (2003); Osterloh and Siewert (2005))
$\tau_{3}$ $=$ $d_{1}-2d_{2}+4d_{3}$
$d_{1}$ $=$ $\psi^{2}_{000}\psi^{2}_{111}+\psi^{2}_{001}\psi^{2}_{110}+\psi^{2}_{010}\psi^{2}_{101}+\psi^{2}_{100}\psi^{2}_{011}$
$d_{2}$ $=$ $\psi_{000}\psi_{111}\psi_{011}\psi_{100}+\psi_{000}\psi_{111}\psi_{101}\psi_{010}$
$+\psi_{000}\psi_{111}\psi_{110}\psi_{001}+\psi_{011}\psi_{100}\psi_{101}\psi_{010}$
$+\psi_{011}\psi_{100}\psi_{110}\psi_{001}+\psi_{101}\psi_{010}\psi_{110}\psi_{001}$
$d_{3}$ $=$ $\psi_{000}\psi_{110}\psi_{101}\psi_{011}+\psi_{111}\psi_{001}\psi_{010}\psi_{100}\ \ ,$
and coincides with the three-qubit hyperdeterminantCayley (1846); Miyake and Wadati (2002). It detects states from the only genuine SL-entangled GHZ-class. $W$-states are not detected; they are instead detected as entangled by the pairwise Concurrence which is distributed along all possible pairs in the state. It is therefore not bipartite and is called genuinely multipartite entangled over the three qubits. However, it is only genuinely pairwise SL-entangled.
Appendix B General properties of optimal decompositions for real states of rank two
@@
$\lambda=\frac{m_{1}}{m_{0}}=\frac{\sqrt{p_{0}(1-p_{0})}\cos\phi_{0}}{\sqrt{p_{1}(1-p_{1})}\cos\phi_{1}}$ (15)
where $m_{i}$ is the weight of the corresponding state. For the mainly interesting case of $\phi_{1}=0$ and $p$ being the probability in the density matrix in consideration, we obtain as a solution to c)
@@
Figure 4: Left: States within the Bloch sphere. For a given line $\overline{x_{1}x_{2}}$ all triangles $z_{0}x_{1}z_{1}$, $z_{0}x_{2}z_{2}$, and $z_{0}x_{2}^{\prime}z_{2}^{\prime}$, are similar. Ratios of their length are determined by the theorem of intersecting lines. Right: The optimal decompositions in the Bloch sphere spanned by GHZGHZ and $W$ state.
@@
B.1 $(1,1)$-decompositions
Here, we collect the formulae obtained for $(1,1)$-decompositions of a pure state at z-coordinate $2p_{1}-1$, where $p_{1}$ corresponds to its probability of the two states $|\psi_{i}\rangle$; $i\in\{N,S\}$. This corresponds to two pure states in the left panel of Fig. 4. We find
@@ B.1 $(1,1)$-decompositions
Next we want to examine whether for a given $(1,1)$ decomposition with $p_{1}$, $p_{2}$, and $p$ it is convenient to split the pure state into two, having a $(1,2)$ decomposition. It can be seen that the ratios $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ are equal with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}_{2}}^{2}}$ (see left panel in Fig. 4). For the new $p^{\prime}_{2}$ of this new state at phase $\phi_{2}$ the calculation gives
@@
Figure 5: Convexification procedure of $|Z3\rangle$ mixed with the corresponding orthogonal state. The curves are made convex with a line touching at $p_{c}=0.0964142$, as is highlighted in the inset, were we plot the difference in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line. The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are shown in blue above both curves.
B.2 $(2,2)$-decompositions
In the following we will assume that two complex conjugated pairs, corresponding to a superposition of two states at $p_{i}$ and angles $\pm\phi_{i}$, $i\in\{1,2\}$, such that both states come to lie inside the Bloch sphere projection. We find
@@ B.2 $(2,2)$-decompositions
Next, for the evaluation of $(2,2)$ from $(2,1)$ decompositions for given $p_{1}$, $\phi_{1}$, and $p$ the ratios of the following quantities are kept fixed (see left panel of Fig. 4): $x_{1}/(2|p_{1}-p|)=x_{2}/(2|p_{2}-p|)=x^{\prime}_{2}/(2|p^{\prime}_{2}-p|)$ with $y^{\prime}_{2}=\pm\sqrt{4p^{\prime}_{2}(1-p^{\prime}_{2})-{x^{\prime}}_{2}^{2}}$. These have to be substituted in $m_{2}(p_{1},\phi_{1};p^{\prime}_{2},\phi^{\prime}_{2})$ where $x^{\prime}_{2}=2\sqrt{p^{\prime}_{2}(1-p^{\prime}_{2})}\cos\phi^{\prime}_{2}$.
Figure 6: Convexification procedure of the two corresponding pure states lying on the line given by the vector $|Z_{1}\rangle+|Z_{2}\rangle$ in a mixed state. The curves are made convex with a line touching at $p_{c}=0.962243$, as is highlighted in the inset, were we plot the difference in $\sqrt{\tau_{3}}$ of the $(1,1)$ decompositions with $|Z_{3}\rangle$ and the convex line. The $(2,1)$ decompositions with $|Z_{3}\rangle$ and $|W\rangle$ are above both curves; they are not shown here.
Appendix C The pure entangled states $|N_{i}\rangle$
We analyze the Bloch sphere of the rank-two mixed state
$\rho[p]=p|GHZ\rangle\langle GHZ|+(1-p)|W\rangle\langle W|$ (33)
One of the pure entangled state is the maximally entangled GHZ stateLohmayer et al. (2006); Viehmann et al. (2012). This is also confirmed through the convexification procedure of $(2,1)$ and $(1,1)$ decompositions which correctly points to the GHZ state. Remains to find the state in one of the three vertical wedges of the Bloch sphere. One possibility is to choose another orthogonal basis for the same Bloch sphere. The one orthogonal decomposition chosen here is the real one of the three pure states of the zero-polytope together with its orthogonal partner.
Figure 7: Left: Difference of $\sqrt{\tau_{3}}$ of $(1,1)$ and $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$. Right: Positions of the entangled pure state in the $(2,1)$ decompositions of $|W\rangle$ and $|Z_{3}\rangle$ for states connecting $|N_{3}\rangle$ and $|M_{3}\rangle$. The inset shows the calculated distance of the plane going through three calculated states to the center of the Bloch sphere.
These are given by
$|Z_{3}\rangle$ $=$ $\sqrt{p_{0}}|GHZ\rangle-\sqrt{1-p_{0}}|W\rangle$ (34)
with $p_{0}=\frac{4\sqrt[3]{2}}{3+4\sqrt[3]{2}}$ Lohmayer et al. (2006). The relevant charcteristic curves for $(1,1)$ decompositions with the W state and $(2,1)$ decompositions with $(|Z_{1}\rangle+|Z_{2}\rangle)/\sqrt{2}$ are shown in Fig. 5 together with their convexification. It shows that at $p_{c}=0.0964142$ the line is a proper convexification, starting from zero and touching the curve for $(2,1)$ decompositions; the $(1,1)$ decompositions are shown in blue. The corresponding pure state is located in $p_{2}(p_{1},\phi_{1},p_{c})=0.00673174$. Since $p_{c}<1/2$, this corresponds to an angle $\theta_{1}=\pi-\arccos{(2p_{c}-1)}=0.164279=9.4125^{\circ}$, which adds up to the angle of $|Z_{3;\perp}\rangle$, which is $\theta_{0}=1.8273=104.697^{\circ}$. The angle of the new state is hence calculated to be $\theta=\theta_{0}+\theta_{1}=1.99158=114.109^{\circ}$. The four tetrahedra are shown in grey in the right panel of Fig. 4 with green edges. This is confirmed from the convexification procedure with the $(1,1)$ decompositions (with $|W\rangle$ here) for eigenstates being situated at an angle $\theta=140^{\circ}$. Invoking the three-fold symmetry for this case leads to the three optimal $(3,1)$ decompositions with the states $|N_{i}\rangle$, $i=1,2,3$, as a tip. This shows as well that in this case $(2,1)$ decompositions are optimal in the direction of $|GHZ\rangle$ and $(1,1)$ decompositions are optimal in the direction of $|W\rangle$. The right panel in Fig 4 shows these $(2,1)$ decompositions as blue lines on the Bloch sphere surface. In Neveling and Osterloh (2024), much effort has been given concerning the distribution of optimal decompositions. These findings hold for even more general SL-invariant tangles $\tau$ than the threetangle, detected by $\sqrt{\tau_{3}}$, and the observations here agree with it. We emphasize that the simplices of $|GHZ\rangle-|Z_{1}\rangle-|Z_{2}\rangle$ is co-planar with $|Z_{1}\rangle-|Z_{2}\rangle-|N_{1}\rangle$; so are the simplices $|Z_{1}\rangle-|N_{1}\rangle-|W\rangle$ and $|Z_{1}\rangle-|N_{3}\rangle-|W\rangle$.
Optimal $(2,1)$ decompositions in between the states $|N_{i}\rangle$ In order to get a glimpse of what is happening in between the lower tetrahedra, we briefly want to see the case in which a density matrix lies within this simplex of $(2,1)$ decompositions. For symmetry reasons, this is the case for the two pure states lying respectively on the vectors corresponding to $|Z_{1}\rangle+|Z_{2}\rangle$ or to $|Z_{3}\rangle+|W\rangle$. The two cases lead to the same result, and we will describe explicitly the first case of $|Z_{1}\rangle+|Z_{2}\rangle$. The corresponding characteristic curves can be viewed in Fig. 6. It is seen that a convexification leads here to a $p_{c}=0.962243$ which corresponds to $p_{2}(p_{0},\pi,p_{c})=0.989858$ for the corresponding pure state on the Bloch sphere surface. This leads to an angle $\theta_{1}=0.201758=11.5599^{\circ}$. As before we have to consider that the two eigenstates lie at an angle $\theta_{0}=2.0539=117.68^{\circ}$. So, the angle of the state $|M_{3}\rangle$ is $\theta=\theta_{0}+\theta_{1}=2.25566=129.24^{\circ}$. The three points corresponding to the states $|M_{i}\rangle$ for $i=1,2,3$ are marked in the right panel in Fig. 1 with white dots. For the further points in the Bloch sphere, we rely on a direct comparison of $(2,1)$ with $(1,1)$ decompositions. We therefore chose the connecting line between $|N_{3}\rangle$ and $|M_{3}\rangle$; the complete solution is obtained by applying the three-fold symmetry. It is important noticing that, at first sight, the effective tangle of $(2,1)$ and $(1,1)$ decompositions yield the same result. That this is not so is seen by looking at the difference between $(1,1)$ and $(2,1)$ decompositions, shown in the left panel of Fig. 7. $\lambda$ refers to the parameter convexly connecting $|N_{3}\rangle$ and $|M_{3}\rangle$, whereas $\lambda_{2}$ is the corresponding parameter for $|W\rangle$ and $|Z_{3}\rangle$. It is seen that the states under consideration get optimised by choosing the $(2,1)$ decomposition. We optimize the tangle with respect to $\lambda_{2}$. The pure states are located in the positions indicated in the right panel of Fig. 7. This leads to the lower blue curves shown in the right panel of Fig. 4 which lie on a circle that is not a grand circle as the ones before; instead it has a distance to the center of the Bloch sphere of $0.0711148$. This is calculated by planes going through three calculated states to the center of the Bloch sphere. In addition to the states $N_{i}$ and $M_{j}$ in between we minimize the tangle on a finite but arbitrary grating. We modified the distance the points should have on the grating. The result did not depend on this choice. The result equals this for the two states $N_{i}$ and the state $M_{j}$ in between. There are numerical errors (due to a finite grating) to this, but they are small (smaller than 44‰) and only occur in the border regions, where the difference to the $(1,1)$ decompositions tends to vanish (see inset of the right panel of Fig. 7). That the states form a circle is confirmed by calculating the derivative with respect to the angle and showing its orthogonality to the original vector (not shown here). The numerical error scales with $\Delta\phi^{2}$ as expected. The normal direction of the circle is (0.57589,0,0.81753)(0.57589,0,-0.81753). Using the symmetry, we get the three points of normal vectors together with a dashed line in blue. The same is done for the azimuthal grand circles.
For the remaining points in the Bloch sphere there is only a single visible state from the zero polytope left which leads to $(1,1)$ decomposition being optimal there. They are depicted in the right panel of Fig. 4 by green lines and red points.